# Formation of clusters and coarsening in weakly interacting diffusions

### v1

@@ Abstract.
This paper studies the clustering behavior of weakly interacting diffusions under the influence of sufficiently localized attractive interaction potentials on the one-dimensional torus. We describe how this clustering behavior is closely related to the presence of discontinuous phase transitions in the mean-field PDE. For local attractive interactions, we employ a new variant of the strict Riesz rearrangement inequality to prove that all global minimizers of the free energy are either uniform or single-cluster states, in the sense that they are symmetrically decreasing.
@@ 1. Introduction
Interacting particle systems and their mean-field limits appear in many applications ranging from physics to mathematical models for swarming, flocking, opinion dynamics, and collective behavior. They are also used in the development of algorithms for sampling and optimization. Often, the emergence of collective behavior, e.g. consensus formation [HK02] or synchronization [Kur81], can be interpreted as a (first- or second-order) phase transition from a disordered to an ordered state; see [CGPS20] and the references therein, in particular [CP10].
@@
In the physical literature, a study of diffusive aggregating particles can be found in [PTA14], where the observed behavior is studied on the PDE level. There, Ostwald ripening is related to volume modes induced by the escape of particles from clusters due to random fluctuations. However, a crucial difference from our study is that the translational modes causing movement and joining of clusters are again deterministically caused by long-range aggregation forces between the clusters, whereas in our model, the interaction potential has finite interaction range and clusters can move only due to the Brownian motions.
@@ 1.1. Setup of the model
If the initial positions $X^{1}_{0},\dots,X^{N}_{0}$ are sampled i.i.d. from some $\rho_{0}\in\mathcal{P}(\mathbb{R}^{d})$ with sufficiently many moments and if $w$ is sufficiently regular (say Lipschitz continuous for instance), then in the limit as the number of particles $N$ goes to infinity, the empirical measure $\mu^{N}_{t}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{X^{i}_{t}}$ at time $t$ converges [Szn91, CD22, CD22a] to $\rho_{t}$, where $(\rho_{t})_{t\geq 0}$ solves the mean-field McKean–Vlasov PDE
@@ 1.2. Summary of the effects
Figure $1.$ Initial clustering for the particle model (left) and for the mean-field PDE (right) for the potential (1.4) with $\ell=0.1$ and $\gamma=10^{4}$. For the particle model, initial positions are evenly spaced on $[0,1)$, that is $X^{i}_{0}=\frac{i}{N}$ for $i=1,\dots,N$. The PDE is started from the state $\frac{100}{98}\chi_{[0.01,0.99]}$ which is a small perturbation of the uniform state. The solution to the PDE is simulated using the Scharfetter–Gummel scheme [SS22].
@@
    Leakage of mass and dynamical metastability: Once the clusters are formed, individual particles can leak from one cluster to another driven by excursions of the Brownian motions which drive the individual particles. This phenomenon is present both in the particle system and mean-field limit. The associated timescale is of the form $e^{\gamma\ell\Delta m}$, where $\Delta=-\inf w$ and $m$ is the cluster’s macroscopic mass, that is the number of particles in the cluster divided by NNFig. 1.3 shows an example for the mass exchange between clusters if the PDE is started initially with three clusters with different masses. Since the timescale for mass leakage $e^{\gamma\ell\Delta m}$ depends exponentially on $m$, the mass exchange among clusters leads to dynamical metastability in the PDE [OR07]: A multi-cluster state can remain stable over a long period of time until it suddenly collapses to a single-cluster state, see Figs. 1.3 and 1.4 for examples.
@@
Figure 1.3. Exchange of mass between three clusters in the mean-field PDE. The potential is given by (1.4) with $\gamma=1000$ and $\ell=0.05$. The initial state is given by a Gaussian mixture (truncated at 0 and 11 and normalized) with centers $(0.1,0.3,0.7)$ and variances $\sigma_{k}^{2}=\frac{\ell}{\gamma m_{k}}$ for $k=1,2,3$ and $m_{1}=0.2,m_{2}=0.3,m_{3}=0.5$. This form of the variances will be explained in more detail Section $3.$
Figure 1.4. Dynamical metastability in the PDE: The PDE is started from a Gaussian mixture with centers $0.25$ and $0.75$ and initial masses $0.45$, $0.55$, respectively. $w$ is given by (1.4) and $\gamma=1000$ and $\ell=0.05$.
In what follows, we will analyse each of these phenomena individually and study the dependence of the associated timescale on the parameters of the system, i.e. the strength of the interaction $\gamma$, the length scale of the interaction $\ell$, and the number of particles in the system $N$. Depending on the relative sizes of the timescales for (E2) and (E3), one of the two effects can dominate the behaviour of the particle system. On the other hand, since the timescale for effect (E2) roughly scales like $N$, the mean-field system does not show effect (E2).
@@ 1.3. Our contributions
  • Generalizing the heuristics by [GPY17] on the initial formation of clusters, we determine the timescale of initial clustering for the mean-field PDE as well as for the particle system under more general initial conditions than [GPY17].
  • We prove for a large family of compactly supported potentials $w$ such as (1.4) that global minimizers of the free energy (1.7) are either the uniform state or single-cluster states in the sense that they are symmetrically decreasing with respect to their center of mass. The latter is true beyond the phase transition, that is for $\gamma>\gamma_{c}$, in which case global minimizers are non-trivial, that is, different from the uniform state. The argument is based on a new variant of the strict Riesz rearrangement inequality which holds true for interaction kernels whose monotony is only locally strict (in contrast to the globally strictly decreasing kernels considered in [Bec93]) and can therefore be applied to certain compactly supported interaction potentials (see also [Sha25, Chapter 5] for related results).
@@ 1.5. Outline
For generic initial conditions and if $\gamma\gg 1$, numerics show that the PDE converges to a single-cluster state. In order to better understand the shape of $\rho_{t}$ as $t\to\infty$, we analyse the minimizers of the free energy (1.7) in Section 3 more closely: First, we prove that for $\ell\ll 1$, the mean-field PDE has a discontinuous phase transition for $\gamma\to\infty$, proving that for $\gamma\gg 1$ large enough, there exist non-trivial global minimizers of the free energy apart from the trivial uniform state $\rho_{\rm unif}\equiv 1$ (see Section 3.1). Moreover, we use a strict version of the Riesz rearrangement inequality to show that all global minimizers of the free energy (1.7) are either uniform or symmetrically decreasing, i.e., single-clusters states, see Section 3.2.
In Section 4 we state and derive a reduced model which is able to describe both the coalescence (E2) and the leakage of mass (E3). We start by recalling the Markovian model from [GPY17] in Section 4.1 and state our extended model with mass exchange in Section 4.2. We show that our new model can explain dynamical metastability of the mean-field limit in Section 4.3 and discuss its limitations in Section 4.4. The heuristic derivation for our model is contained in Section 4.5. As a byproduct, we obtain also a jump process description in Section 4.6.
We conclude the paper by discussing possible future directions in Section 5. Moreover, in Appendix A, we formulate a conjecture on the convergence of (1.1) to the Modified Massive Arratia flow [Kon17, KR19] in a suitable scaling limit.
@@ 2.1. Background: Initial clustering after [GPY17] via linearization
Linearizing the PDE (1.6) around the uniform equilibrium state $\rho_{\rm unif}(x)\equiv 1$ for a small perturbation ${\rho}_{\rm pert}$, we obtain the linear PDE
@@
$\psi(k)$ $:=\operatorname{Re}\left(\rho_{\rm unif}k\int_{[0,1]}W^{\prime}_{\gamma,\ell}(x)e^{-ikx}\mathop{}\!\mathup{d}x-k^{2}\right)=-k^{2}\left(\hat{W}_{\gamma,\ell}(k)+1\right),$ (2.3)
@@
see also [CP10], where $\gamma_{\sharp}>0$ is true for sufficiently small $\ell>0$ since for those $\ell$ it holds that $\widehat{W}_{1,\ell}(1)=\ell\int_{\operatorname{supp}w}w(x)\cos(\ell x)\mathop{}\!\mathup{d}x<0$.
@@
On the other hand, if $\gamma>\gamma_{\sharp}$, then $\max_{k\in\mathcal{K}}\psi(k)>0$, the system (2.1) is unstable and clustering occurs by means of the dominant Fourier mode $k_{\rm max}$ with maximal growth rate. The estimated number of initial clusters is determined by $k_{\rm max}$ and given by
$n_{\rm clusters}=\frac{k_{\rm max}}{2\pi}\in\mathbb{N}.$ (2.7)
In particular, the inter-cluster distance is given by $d_{\rm clusters}=1/n_{\rm clusters}=2\pi/k_{\rm max}$. Clustering in the linearized PDE (2.1) with a number of $n_{\rm clusters}$ occurs whenever the initial condition ${\rho}_{\rm pert}(0,{\color[rgb]{.75,.75,.75}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.75,.75}\pgfsys@color@gray@stroke{.75}\pgfsys@color@gray@fill{.75}{\bullet}})$ is sufficiently close to a constant function and ${\hat{\rho}}_{\rm pert}(0,k_{\rm max})=\overline{{\hat{\rho}}_{\rm pert}(0,-k_{\rm max})}\neq 0$. Let us assume such an initial condition for (2.1). Then, since the dominant modes ${\hat{\rho}}_{\rm pert}(t,\pm k_{\rm max})$ grow exponentially as
$\lvert{\hat{\rho}}_{\rm pert}(t,\pm k_{\rm max})\rvert=e^{\psi(\pm k_{\rm max})t}\lvert{\hat{\rho}}_{\rm pert}(0,\pm k_{\rm max})\rvert\qquad\text{for $t>0$},$ (2.8)
the initial clustering time for (2.1) is of the order
@@
As we will argue in the following, initial clustering in the regime $\gamma>\gamma_{\sharp}$ occurs both for the PDE (1.6) when started from a perturbation of the uniform state which has a non-zero $k_{\rm max}$-Fourier mode and for the particle system.
@@ 2.2. Initial clustering in the PDE
The previous section shows that for $\gamma>\gamma_{\sharp}$, starting the PDE from a small perturbation $\rho_{0}\in\mathcal{P}([0,1])\cap L^{2}([0,1])$ of the uniform state $\rho_{\rm unif}\equiv 1$ such that ${\hat{\rho}}_{\rm pert}(0,k_{\rm max})\neq 0$ yields initial clustering with the same number of clusters $n_{\rm clusters}$ as the particle system. This is also observed numerically, see (Fig. 1.1, right). The initial clustering timescale is then of the order
@@ 2.3. Initial clustering in the particle system
converge to a Gaussian-valued measure $\mu$, that is, for each $h\in C_{b}([0,1])$, the real-valued random variable $\int_{[0,1]}h(x)\mu({\rm d}x)$ is $\mathcal{N}(0,\sigma^{2}(h))$-distributed, where $\sigma^{2}(h)=\int\bigl(h(x)-\int h(y)\rho_{0}({\rm d}y)\bigr)^{2}\rho_{0}({\rm d}x)$. Writing the initial empirical measure $\mu^{N}_{0}=\frac{1}{N}\sum_{j=1}^{N}\delta_{X^{j}_{0}}$ in the form
$\mu^{N}_{0}\approx\rho_{\rm unif}+(\rho_{0}-\rho_{\rm unif})+\frac{\mu}{\sqrt{N}},$ (2.12)

### v2

@@ Abstract.
This paper studies the clustering behaviour of weakly interacting diffusions under the influence of sufficiently localized attractive interaction potentials on the one-dimensional torus. We describe how this clustering behaviour is closely related to the presence of discontinuous phase transitions in the mean-field PDE. For local attractive interactions, we employ a new variant of the strict Riesz rearrangement inequality to prove that all global minimisers of the free energy are either uniform or single-cluster states, in the sense that they are symmetrically decreasing.
@@ 1. Introduction
Interacting particle systems and their mean-field limits appear in many applications ranging from physics to mathematical models for swarming, flocking, opinion dynamics, and collective behaviour. They are also used in the development of algorithms for sampling and optimization. Often, the emergence of collective behaviour, e.g. consensus formation [HK02] or synchronization [Kur81], can be interpreted as a (first- or second-order) phase transition from a disordered to an ordered state; see [CGPS20] and the references therein, in particular [CP10].
@@
In the physical literature, a study of diffusive aggregating particles can be found in [PTA14], where the observed behaviour is studied on the PDE level. There, Ostwald ripening is related to volume modes induced by the escape of particles from clusters due to random fluctuations. However, a crucial difference from our study is that the translational modes causing movement and joining of clusters are again deterministically caused by long-range aggregation forces between the clusters, whereas in our model, the interaction potential has finite interaction range and clusters can move only due to the Brownian motions.
@@ 1.1. Setup of the model
If the initial positions $X^{1}_{0},\dots,X^{N}_{0}$ are sampled i.i.d. from some $\rho_{0}\in\mathcal{P}(\mathbb{R}^{d})$ with sufficiently many moments and if $w$ is sufficiently regular, then in the limit as the number of particles $N$ goes to infinity, the empirical measure $\mu^{N}_{t}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{X^{i}_{t}}$ at time $t$ converges [Szn91, CD22, CD22a] to $\rho_{t}$, where $(\rho_{t})_{t\geq 0}$ solves the mean-field McKean–Vlasov PDE
@@ 1.2. Summary of the effects
Figure $1.$ Initial clustering for the particle model (left) and for the mean-field PDE (right) for the potential (1.4) with $\ell=0.1$ and $\gamma=10^{4}$. For the particle model, initial positions are evenly spaced on $[0,1)$, that is $X^{i}_{0}=\frac{i}{N}$ for $i=1,\dots,N$. The PDE is started from the state $\frac{100}{98}\chi_{[0.01,0.99]}$ which is a small perturbation of the uniform state. The solution to the PDE is simulated using the Scharfetter–Gummel scheme [SS22], please see https://gitlab.uni-ulm.de/clustercoarsening/cluster-coarsening-paper for the code repository used to generate the figures in this paper.
@@
    Leakage of mass and dynamical metastability: Once the clusters are formed, individual particles can leak from one cluster to another driven by excursions of the Brownian motions which drive the individual particles. This phenomenon is present both in the particle system and mean-field limit. The associated timescale is of the form $e^{\gamma\ell\Delta m}$, where $\Delta=-\inf w$ and $m$ is the cluster’s macroscopic mass, that is the number of particles in the cluster divided by $N$. In the case of the microscopic model, (E3) competes with (E2) depending on the relative sizes of the timescales of the two phenomena, which are of the order $e^{\gamma\ell\Delta m}$ and $N$, respectively.  Fig. 1.3 shows an example for the mass exchange between clusters if the PDE is started initially with three clusters with different masses. Since the timescale for mass leakage $e^{\gamma\ell\Delta m}$ depends exponentially on $m$, the mass exchange among clusters leads to dynamical metastability in the PDE [OR07]: A multi-cluster state can remain stable over a long period of time until it suddenly collapses to a single-cluster state, see Figs. 1.3 and 1.4 for examples.
@@
Figure 1.3. Exchange of mass between three clusters in the mean-field PDE. The potential is given by (1.4) with $\gamma=1000$ and $\ell=0.05$. The initial state is given by a Gaussian mixture (truncated at 0 and 11 and normalized) with centres $(0.1,0.3,0.7)$ and variances $\sigma_{k}^{2}=\frac{\ell}{\gamma m_{k}}$ for $k=1,2,3$ and $m_{1}=0.2,m_{2}=0.3,m_{3}=0.5$. This form of the variances will be explained in more detail in Section $3.$
Figure 1.4. Dynamical metastability in the PDE: The PDE is started from a Gaussian mixture with centres $0.25$ and $0.75$ and initial masses $0.45$, $0.55$, respectively. $w$ is given by (1.4) and $\gamma=1000$ and $\ell=0.05$.
In what follows, we will analyse each of these phenomena individually and study the dependence of the associated timescale on the parameters of the system, i.e. the strength of the interaction $\gamma$, the length scale of the interaction $\ell$, and the number of particles in the system $N$. As already mentioned above, depending on the relative sizes of the timescales for (E2) and (E3), one of the two effects can dominate the behaviour of the particle system. On the other hand, since the timescale for effect (E2) roughly scales like $N$, the mean-field system does not show effect (E2).
@@ 1.3. Our contributions
  • Generalising the heuristics by [GPY17] on the initial formation of clusters, we determine the timescale of initial clustering for the mean-field PDE as well as for the particle system under more general initial conditions than [GPY17]. Moreover, we determine the point of linear stability $\gamma_{\sharp}$ explicitly.
  • We prove for a large family of compactly supported potentials $w$ such as (1.4) that global minimisers of the free energy (1.7) are either the uniform state or single-cluster states in the sense that they are symmetrically decreasing with respect to their centre of mass. The latter is true beyond the phase transition, that is, for $\gamma>\gamma_{c}$, in which case global minimisers are non-trivial, that is, different from the uniform state. The argument is based on a new variant of the strict Riesz rearrangement inequality which holds true for interaction kernels whose monotony is only locally strict (in contrast to the globally strictly decreasing kernels considered in [Bec93]) and can therefore be applied to certain compactly supported interaction potentials (see also [Sha25, Chapter 5] for related results).
@@ 1.5. Outline
For generic initial conditions and if $\gamma\gg 1$, numerics show that the PDE converges to a single-cluster state. In order to better understand the shape of $\rho_{t}$ as $t\to\infty$, we analyse the minimisers of the free energy (1.7) in Section 3 more closely: First, we prove that for $\ell\ll 1$, the mean-field PDE has a discontinuous phase transition for $\gamma\to\infty$, proving that for $\gamma\gg 1$ large enough, there exist non-trivial global minimisers of the free energy apart from the trivial uniform state $\rho_{\rm unif}\equiv 1$ (see Section 3.1). Moreover, we use a strict version of the Riesz rearrangement inequality to show that all global minimisers of the free energy (1.7) are either uniform or symmetrically decreasing, i.e., single-clusters states, see Section 3.2.
In Section 4 we state and derive a reduced model which is able to describe both the coalescence (E2) and the leakage of mass (E3). We start by recalling the Markovian model from [GPY17] in Section 4.1 and state our extended model with mass exchange in Section 4.2. We show that our new model can explain dynamical metastability of the mean-field limit in Section 4.3 and discuss its limitations in Section 4.4. The heuristic derivation for our model is contained in Section 4.5. As a byproduct, we also obtain a jump process description in Section 4.6.
We conclude the paper by discussing possible future directions in Section 5. Moreover, in Appendix A, we formulate a conjecture on the convergence of (1.1) to the Modified Massive Arratia flow [Kon17, KR19] in a suitable scaling limit. Finally, Appendix B provides a proof for the computation of the mass leakage times used in our extended model.
@@ 2.1. Background: Initial clustering after [GPY17] via linearization
Assuming that $\rho_{t}=\rho_{\rm unif}+{\rho}_{\rm pert}$, where $\rho_{\rm unif}(x)\equiv 1$ is the uniform uniform equilibrium state and ${\rho}_{\rm pert}$ is a small perturbation, we linearize the McKean–Vlasov PDE (1.6) around $\rho_{\rm unif}$ to obtain the linear PDE
@@
$\psi(k)$ $:=\operatorname{Re}\left(i\rho_{\rm unif}k\int_{[0,1]}W^{\prime}_{\gamma,\ell}(x)e^{-ikx}\mathop{}\!\mathup{d}x-k^{2}\right)=-k^{2}\left(\hat{W}_{\gamma,\ell}(k)+1\right),$ (2.3)
@@
see also [CP10], where $\gamma_{\sharp}>0$ is true for sufficiently small $\ell>0$ since for those $\ell$ it holds that $\widehat{W}_{1,\ell}(2\pi)=\ell\int_{\operatorname{supp}w}w(x)\cos(2\pi\ell x)\mathop{}\!\mathup{d}x<0$. Below in Lemma 2.1, we will show that the minimum in (2.6) is always attained at the first Fourier mode $k=\pm 2\pi$.
@@
On the other hand, if $\gamma>\gamma_{\sharp}$, then $\max_{k\in\mathcal{K}}\psi(k)>0$ and the system (2.1) is unstable. Therefore, provided that the mode $k_{\rm max}$ with maximal growth rate is already present initially, that is, ${\hat{\rho}}_{\rm pert}(0,k_{\rm max})=\overline{{\hat{\rho}}_{\rm pert}(0,-k_{\rm max})}\neq 0$, the short-time behaviour of (2.1) will be dominated by $e^{\pm ik_{\max}t}$, whose amplitude grows according to
$\lvert{\hat{\rho}}_{\rm pert}(t,\pm k_{\rm max})\rvert=e^{\psi(\pm k_{\rm max})t}\lvert{\hat{\rho}}_{\rm pert}(0,\pm k_{\rm max})\rvert\qquad\text{for $t>0$}.$ (2.7)
In particular, (2.1) exhibits initial clustering by means of $e^{\pm ik_{\max}t}$ on a timescale of order
@@
The estimated number of initial clusters is determined by $k_{\rm max}$ and given by
$n_{\rm clusters}=\frac{k_{\rm max}}{2\pi}\in\mathbb{N},$ (2.9)
with the inter-cluster distance being $d_{\rm clusters}=1/n_{\rm clusters}=2\pi/k_{\rm max}$. We will now describe the occurrence and the timescale for initial clustering for both the PDE (1.6) and the particle system.
@@ 2.2. Initial clustering in the PDE
Assume that $\gamma>\gamma_{\sharp}$ and that the PDE is started from a small perturbation $\rho_{0}\in\mathcal{P}([0,1])\cap L^{2}([0,1])$ of the uniform state $\rho_{\rm unif}\equiv 1$ such that $\widehat{\rho_{0}}(0,k_{\rm max})\neq 0$. Then, the short-time behaviour of the PDE is governed by its linearization (2.1) around $\rho_{\rm unif}$ and the previous section shows that the PDE exhibits initial clustering with a number of $n_{\rm clusters}$ clusters. This is also observed numerically, see (Fig. 1.1, right). Because of (2.8), the initial clustering timescale is then of the order
@@ 2.3. Initial clustering in the particle system
converge to a Gaussian-valued measure $\mu$, that is, for each $h\in C_{b}([0,1])$, the real-valued random variable $\int_{[0,1]}h(x)\mu({\rm d}x)$ is $\mathcal{N}(0,\sigma^{2}(h))$-distributed, where $\sigma^{2}(h)=\int\bigl(h(x)-\int h(y)\rho_{0}({\rm d}y)\bigr)^{2}\rho_{0}({\rm d}x)$. Writing the initial empirical measure $\mu^{N}_{0}$ in the form
$\mu^{N}_{0}=\frac{1}{N}\sum_{j=1}^{N}\delta_{X^{j}_{0}}\approx\rho_{\rm unif}+(\rho_{0}-\rho_{\rm unif})+\frac{\mu}{\sqrt{N}},$ (2.12)
@@
2.4. Calculating the point of linear stability $\gamma_{\sharp}$
Next, we prove that the minimum in (2.6) is always attained at the first Fourier mode. This gives an explicit expression for $\gamma_{\sharp}$, showing that $\gamma_{\sharp}\asymp\ell^{-2}$ as $\ell\downarrow 0$.
Lemma 2.1.
Under 1.1, the point of linear stability $\gamma_{\sharp}$ as defined in (2.6) equals
$\gamma_{\sharp}=-\frac{1}{\widehat{W}_{\gamma=1,\ell}(2\pi)}$ (2.15)
whenever $0<\ell<\frac{1}{4s_{w}}$, where $\operatorname{supp}w=[-s_{w},s_{w}]$. Moreover, we have the asymptotic equivalence
$\gamma_{\sharp}\sim\frac{1}{m_{w}}\ell^{-2}\qquad\text{as $\ell\downarrow 0$},\qquad\text{where $m_{w}=-\int_{\mathbb{R}}w(x)\mathop{}\!\mathup{d}x>0$}.$ (2.16)
Proof.
(2.15) follows once we have proven that $\widehat{W}_{\gamma=1,\ell}(k)$ is minimized on $(2\pi\mathbb{Z})\setminus\{0\}$ at $k=\pm 2\pi$. To show this, fix $k\in 2\pi\mathbb{N}_{+}$. We decompose the $k$-th Fourier mode of $W_{\gamma=1,\ell}$ as follows,
$\widehat{W}_{\gamma=1,\ell}(k)$ $=\ell\int_{-\frac{1}{2}}^{\frac{1}{2}}w\left(\frac{x}{\ell}\right)\cos(kx)\mathop{}\!\mathup{d}x=2\ell\int_{0}^{\infty}w\left(\frac{x}{\ell}\right)\cos(kx)\mathop{}\!\mathup{d}x$ (2.17)
$=2\ell\int_{0}^{\frac{\pi}{2k}}w\left(\frac{x}{\ell}\right)\cos(kx)\mathop{}\!\mathup{d}x+2\ell\sum_{m=0}^{\infty}\int_{\frac{\pi}{2k}(1+4m)}^{\frac{\pi}{2k}(3+4m)}w\left(\frac{x}{\ell}\right)\cos(kx)\mathop{}\!\mathup{d}x$ (2.18)
$\qquad\qquad+2\ell\sum_{m=0}^{\infty}\int_{\frac{\pi}{2k}(3+4m)}^{\frac{\pi}{2k}(5+4m)}w\left(\frac{x}{\ell}\right)\cos(kx)\mathop{}\!\mathup{d}x,$ (2.19)
where we split the integrals at the zeros of $\cos(kx)$. Since $w$ has compact support, the sums are finite. Using the substitution $x\mapsto x+\frac{\pi}{k}$, we have for each $m\in\mathbb{N}$ that
$\int_{\frac{\pi}{2k}(1+4m)}^{\frac{\pi}{2k}(3+4m)}w\left(\frac{x}{\ell}\right)\cos(kx)\mathop{}\!\mathup{d}x+\int_{\frac{\pi}{2k}(3+4m)}^{\frac{\pi}{2k}(5+4m)}w\left(\frac{x}{\ell}\right)\cos(kx)\mathop{}\!\mathup{d}x$ (2.20)
$={}$ $\int_{\frac{\pi}{2k}(1+4m)}^{\frac{\pi}{2k}(3+4m)}w\left(\frac{x}{\ell}\right)\cos(kx)+w\left(\frac{x}{\ell}+\frac{\pi}{\ell k}\right)\cos\left(kx+\pi\right)\mathop{}\!\mathup{d}x$ (2.21)
$={}$ $\int_{\frac{\pi}{2k}(1+4m)}^{\frac{\pi}{2k}(3+4m)}\left(w\left(\frac{x}{\ell}\right)-w\left(\frac{x}{\ell}+\frac{\pi}{\ell k}\right)\right)\cos(kx)\mathop{}\!\mathup{d}x.$ (2.22)
Since $w$ is non-decreasing on $0,\infty)$, (2.22) is non-negative and we obtain
$\widehat{W}_{\gamma=1,\ell}(k)$ $\geq\ell\int_{-\frac{\pi}{2k}}^{\frac{\pi}{2k}}w\left(\frac{x}{\ell}\right)\cos(kx)\mathop{}\!\mathup{d}x.$ (2.23)
Moreover, since $\cos(kx)\leq\cos(2\pi x)$ for all $x\in[0,\frac{\pi}{2k}]$, we have that
$\widehat{W}_{\gamma=1,\ell}(k)$ $\geq\ell\int_{-\frac{\pi}{2k}}^{\frac{\pi}{2k}}w\left(\frac{x}{\ell}\right)\cos(kx)\mathop{}\!\mathup{d}x>\ell\int_{-\frac{\pi}{2k}}^{\frac{\pi}{2k}}w\left(\frac{x}{\ell}\right)\cos(2\pi x)\mathop{}\!\mathup{d}x$ (2.24)
$\geq\ell\int_{-\frac{1}{4}}^{\frac{1}{4}}w\left(\frac{x}{\ell}\right)\cos(2\pi x)\mathop{}\!\mathup{d}x.$ (2.25)
The inequality in the second step is strict since $w$ is negative on $(-s_{w},s_{w})$ and $\cos(kx)<\cos(2\pi x)$ for all $x\in(0,\frac{\pi}{2k}]$. Assuming that $0<\ell<\frac{1}{4s_{w}}$, we have that
$\ell\int_{-\frac{1}{4}}^{\frac{1}{4}}w\left(\frac{x}{\ell}\right)\cos(2\pi x)\mathop{}\!\mathup{d}x=\ell\int_{-\frac{1}{2}}^{\frac{1}{2}}w\left(\frac{x}{\ell}\right)\cos(2\pi x)\mathop{}\!\mathup{d}x=\widehat{W}_{\gamma=1,\ell}(2\pi).$ (2.26)
In particular, $\widehat{W}_{\gamma=1,\ell}(k)$ is minimized on $(2\pi\mathbb{Z})\setminus\{0\}$ at $k=\pm 2\pi$.
The asymptotic behaviour (2.16) follows from the fact that $\ell^{-1}w(\ell^{-1}{\color[rgb]{.75,.75,.75}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.75,.75}\pgfsys@color@gray@stroke{.75}\pgfsys@color@gray@fill{.75}{\bullet}})\to-m_{w}\delta_{0}$ in the sense of distributions as $\ell\downarrow 0$. In fact, we have that
$\ell^{-2}\widehat{W}_{\gamma=1,\ell}(2\pi)=\int_{\mathbb{R}}w\left(x\right)\cos(2\pi\ell x)\mathop{}\!\mathup{d}x\to\int_{\mathbb{R}}w(x)\mathop{}\!\mathup{d}x\qquad\text{as $\ell\downarrow 0$}.$ (2.27)